Temperature changes in seawater affect the
speed at which sound moves through it. A wave
with a length of 2.00 m, travelling at a speed of
1500 m/s, reaches a section of warm water where
the speed is 1550 m/s. What would you expect
the wavelength in the warmer water to be?
1
Expert's answer
2020-06-22T11:08:06-0400
"f_1=f_2\\\frac{v_1}{\lambda_1}=\frac{v_2}{\lambda_2}\\\frac{1500}{2}=\frac{1550}{\lambda_2}"

"\lambda_2=2.07\ m"